You are breathing at a rate of 20 breaths per minute in a warm, dry environment.
Each time you take a breath, you breathe in 0.5×10−3 kg of air at a temperature of 24°C and with a relative humidity of 20%. You breathe out the same mass of air but at a temperature of 36°C and a relative humidity of 100%.
How much water vapour do you exhale in one hour under these circumstances? (to 2 s.f and in g)
1
Expert's answer
2020-05-12T10:23:32-0400

Solution.

"n=" 20 breaths per minute;

"m_a=0.5\u00d710^{\u22123} kg;"

"t_1=24\u00b0C;"

"\phi_1=20\%;"

"t_2=36\u00b0C;"

"\phi_2=100\%;"

"\tau=1 hour;"

You take 20×60=1200 breaths per hour

In an hour you inhale 1200×0.5×10−3 kg=600×10−3 kg;

"\phi_1=\dfrac{\rho_{1a}}{\rho_{1s}}\sdot100\%;\implies \rho_{1a}=\dfrac{\phi_1\sdot\rho_{1s}}{100\%};"

"\rho_{1s}=21.8g/m^3" - saturated vapor density at a temperature of 24oC (table);

"\rho_{1a}=\dfrac{20\%\sdot21.8g/m^3}{100\%}=4.36g/m^3;"

"\phi_2=\dfrac{\rho_{2a}}{\rho_{2s}}\sdot100\%;\implies \rho_{2a}=\dfrac{\phi_2\sdot\rho_{2s}}{100\%};"

"\rho_{2s}=42.1g/m^3;"

"\rho_{2a}=42.1g/m^3;"

The density of dry air is equal to 1.189kg/m3;

"m_a=\rho V\implies V=\dfrac{m_a}{\rho}";

"V=\dfrac{600\sdot 10^{-3}kg}{1.189kg/m^3}=504.63\sdot10^{-3}m^3;"

"m_v=\rho_{2a}V;"

"m_v=42.1g/m^3\sdot504.63\sdot10^{-3}m^3=21.24g;"

Answer: "m_v=21.24g."